# Towards Intelligible Human-Robot Interaction: An Active Inference Approach to Occluded Pedestrian Scenarios

### v1

@@ Appendix D Details of Comparison with Baseline Methods
PPO-LSTM $0.650\pm 0.044$ $0.350\pm 0.044$ $\mathbf{4.172\pm 0.005}$ $\mathbf{3.772\pm 0.087}$ $\mathbf{8.493\pm 0.400}$
Ours $\mathbf{0.925\pm 0.024}$ $\mathbf{0.075\pm 0.024}$ $8.172\pm 0.076$ $3.651\pm 0.093$ $1.519\pm 0.081$
@@
PPO-LSTM $0.733\pm 0.040$ $0.267\pm 0.040$ $\mathbf{4.178\pm 0.005}$ $3.932\pm 0.058$ $\mathbf{10.000\pm 0.000}$
Ours $\mathbf{0.850\pm 0.033}$ $\mathbf{0.150\pm 0.033}$ $8.692\pm 0.125$ $\mathbf{4.249\pm 0.093}$ $1.637\pm 0.065$
@@
PPO-LSTM $0.833\pm 0.034$ $0.167\pm 0.034$ $\mathbf{4.195\pm 0.002}$ $\mathbf{3.616\pm 0.049}$ $\mathbf{9.708\pm 0.166}$
Ours $\mathbf{0.992\pm 0.008}$ $\mathbf{0.008\pm 0.008}$ $8.336\pm 0.051$ $3.267\pm 0.023$ $1.056\pm 0.044$
@@
PPO-LSTM $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $\mathbf{4.200\pm 0.000}$ $\mathbf{4.947\pm 0.048}$ $\mathbf{10.000\pm 0.000}$
Ours $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $9.832\pm 0.129$ $3.661\pm 0.051$ $1.475\pm 0.058$
@@
PPO-LSTM $0.408\pm 0.045$ $0.592\pm 0.045$ $\mathbf{4.188\pm 0.005}$ $3.659\pm 0.077$ $\mathbf{10.000\pm 0.000}$
Ours $\mathbf{0.967\pm 0.016}$ $\mathbf{0.033\pm 0.016}$ $8.186\pm 0.094$ $\mathbf{4.678\pm 0.107}$ $2.039\pm 0.055$

### v2

@@ Appendix D Details of Comparison with Baseline Methods
PPO-LSTM $0.650\pm 0.044$ $0.350\pm 0.044$ $\textbf{4.172}\pm\textbf{0.005}$ $\textbf{3.772}\pm\textbf{0.087}$ $\textbf{8.493}\pm\textbf{0.400}$
Ours $\textbf{0.925}\pm\textbf{0.024}$ $\textbf{0.075}\pm\textbf{0.024}$ $8.172\pm 0.076$ $3.651\pm 0.093$ $1.519\pm 0.081$
@@
PPO-LSTM $0.733\pm 0.040$ $0.267\pm 0.040$ $\textbf{4.178}\pm\textbf{0.005}$ $3.932\pm 0.058$ $\textbf{10.000}\pm\textbf{0.000}$
Ours $\textbf{0.850}\pm\textbf{0.033}$ $\textbf{0.150}\pm\textbf{0.033}$ $8.692\pm 0.125$ $\textbf{4.249}\pm\textbf{0.093}$ $1.637\pm 0.065$
@@
PPO-LSTM $0.833\pm 0.034$ $0.167\pm 0.034$ $\textbf{4.195}\pm\textbf{0.002}$ $\textbf{3.616}\pm\textbf{0.049}$ $\textbf{9.708}\pm\textbf{0.166}$
Ours $\textbf{0.992}\pm\textbf{0.008}$ $\textbf{0.008}\pm\textbf{0.008}$ $8.336\pm 0.051$ $3.267\pm 0.023$ $1.056\pm 0.044$
@@
PPO-LSTM $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $\textbf{4.200}\pm\textbf{0.000}$ $\textbf{4.947}\pm\textbf{0.048}$ $\textbf{10.000}\pm\textbf{0.000}$
Ours $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $9.832\pm 0.129$ $3.661\pm 0.051$ $1.475\pm 0.058$
@@
PPO-LSTM $0.408\pm 0.045$ $0.592\pm 0.045$ $\textbf{4.188}\pm\textbf{0.005}$ $3.659\pm 0.077$ $\textbf{10.000}\pm\textbf{0.000}$
Ours $\textbf{0.967}\pm\textbf{0.016}$ $\textbf{0.033}\pm\textbf{0.016}$ $8.186\pm 0.094$ $\textbf{4.678}\pm\textbf{0.107}$ $2.039\pm 0.055$